# Learning Shared Representations from Unpaired Data

### v1

@@ A.7 Qualitative Results
A.8 Extended Ablation Study
@@
Table 8: MMD Retrieval results. R@10 retrieval results on vision-language ([Uncaptioned image]) and vision-vision ([Uncaptioned image]) datasets from each modality to another: image-to-text (I2T), text-to-image (T2I), edges-to-shoes (E2S), shoes-to-edges (S2E), note-to-measurement (N2M), measurement-to-note (M2N). SUE results are significantly better, supporting the key role of SE in the universal embedding concept and SUE, as well as the necessity of the CCA step in the current implementation.
@@ A.9 Ablation Visualization
Appendix B Modality Gap
Recent research involved the modality gap in multimodal learning Liang et al. (2022); Jiang et al. (2023); Schrodi et al. (2025). That is, the embeddings of the two modalities of contrastive-based methods are located in two completely separate regions of the embedding space. Some efforts have been made to mitigate this gap Jiang et al. (2024); Eslami and de Melo (2025). In particular, the modality gap means that the embeddings are not universal, as the modalities lay on different intrinsic manifolds. In contrast, a universal embedding should have no modality gap.
As visualized in Fig. 14, SUE indeed does not contain a modality gap. The embedding discovered by SUE is a shared manifold of the two modalities (unlike CLIP, for instance).
@@ Appendix C Relative CLIP Score
The qualitative examples in Fig. 2 and Fig. 4a depict the retrieval weakness in reliably measuring semantic coherence across modalities. We address this by presenting a soft variant of Recall@$k$ for the text-image domain, which we name Relative CLIP Score (RCS). CLIP score (CS) is a popular measure of similarity between images Hessel et al. (2021), but is difficult to interpret. Following this idea, we define the RCS between a pair of an image query and a retrieved text as the CS between them, relative to the CS between the image and its true pair (the RCS between a text query and a retrieved image is defined analogously). This results in a measurement that spans (approximately) between 0 to 1.
@@
Appendix D Technical Details
D.1 Datasets
Table 10: Datasets’ properties. #unpaired indicates the number of unpaired samples used by SUE; #paired indicates the number of paired samples used by SUE. Model 1 refers to the unimodal encoder for the first modality, and Model 2 corresponds to the second modality. More details on the Unimodal encoders can be found in App. D.2
. Dataset Modality 1 Modality 2 Model 1 Model 2 #unpaired #paired Flickr30k Images Captions DINOv2 SBERT 27794 500 Polyvore Images Captions FashionSigLIP SBERT 17190 500 MSCOCO Images Captions DINOv2 SBERT 8550 100 Edges2Shoes Images Edges (images) DINOv2 VICReg 12690 50 caption-FFHQ Images Captions DINOv2 SBERT 16958 1000 Office31 Images Images VICReg DINOv2 8000 1000
@@ D.1 Datasets
D.2 Unimodal Models
As extensively discussed in Sec. 1 and Sec. 4, our first step is extracting meaningful (unimodal) embeddings that capture semantic similarity, using pre-trained unimodal models. The unimodal models used for the different datasets include: DINOv2 (ViT-B/14 distilled) Oquab et al. (2023), VICReg Bardes et al. (2021), SBERT (all-MiniLM-L6-v2) Reimers and Gurevych (2019), he Marqo-FashionSigLIP image encoder [^2] https://www.marqo.ai/blog/search-model-for-fashion, and the MPMTS Measurements encoder King et al. (2023). Tab. 10 details the models used for each dataset. Tab. 11 details the key properties of each unimodal encoder.
@@ D.2 Unimodal Models
D.3 Generation and zero-shot
@@
After obtaining the universal image embeddings from the caption-FFHQ dataset using SUE, we encoded the entire dataset of images with a pre-trained GAN inversion model Wang et al. (2022) to obtain the latent representation for all images. Next, we trained a converter model to map the universal image embedding of an input image to its corresponding latent. To achieve this, we used an MSE loss to align the converter’s output with the latent obtained from the GAN’s encoder.
@@
For this experiment, we collected random images of 7 object categories from the ImageNet dataset Deng et al. (2009): cat, dog, man, woman, mountain, shore, snow. Each label was converted into the following sentence: "A photo of a []". For each input image, we computed its universal embedding using SUE, which was trained on Flickr30k and then calculated the cosine similarity between the image embedding and the universal embeddings of all label texts. The label for the image was predicted by selecting the text label that had the highest cosine similarity to the image’s universal embedding.
D.4 Cross-domain Classification
@@
For unimodal models, we used VICReg Bardes et al. (2021) encoder for the Amazon domain and DINOv2 Oquab et al. (2023) for the DSLR domain. After obtaining universal embeddings from SUE on the test set, we constructed a KNN classifier using 5 neighbors on one domain and evaluated it on the other, as explained in App. A.3.
D.5 Graph Construction
@@
As unimodal models are often trained using cosine similarity (e.g., Reimers and Gurevych (2019); Oquab et al. (2023)), we analogously defined $d$ to be the cosine distance:
@@
D.6 Training procedure
For the parametric computation of the SE, we adopted the training process presented in Shaham et al. (2018), and trained a neural network capable of approximating the SE on the training set. Each modality was trained independently, and upon completion, the resulting parametric maps were used to compute the SE for the test points of each modality. For further details on the training process for this parametric map, we refer the reader to Shaham et al. (2018).
@@ D.6 Training procedure
D.7 Data split and preparation
@@ D.8 Hyper-parameters
For the numeric SE, we constructed the graph as outlined in Sec. D.5, using $k=100$ for each point. The Laplacian matrix used is the random walk Laplacian, and we selected 10 eigenvectors, which correspond to the output dimension of the SE.
@@
Computing the SE using the parametric approach involves training a neural network. The network architecture consists of an MLP with hidden layers of sizes $4096$, $4096$, and $1024$ for both networks (one per modality). We set the batch size to 4096, and the learning rate to $10^{-4}$ with a decay factor of 0.1, and the training was run for 100 epochs. The optimizer used is Adam, and the learning rate is adjusted using the PyTorch ReduceLROnPlateau scheduler with a patience of 10. We used the same graph construction as used for the numeric SE, outlined in Sec. D.5.
@@
To train the residual network, we employed an MLP architecture with hidden layers of size $128$, $128$, and $128$, incorporating a residual connection from the input to the output. The optimizer used is AdamW, with a learning rate set to $10^{-3}$. The network was trained for 100 epochs.

### v2

@@ A.9 Qualitative Results
A.10 Extended Ablation Study
@@
Table 9: Extended Ablation. R@10 retrieval results on vision-language ([Uncaptioned image]) and vision-vision ([Uncaptioned image]) datasets from each modality to another: image-to-text (I2T), text-to-image (T2I), edges-to-shoes (E2S), shoes-to-edges (S2E), note-to-measurement (N2M), measurement-to-note (M2N). SUE results are significantly better, supporting the key role of SE in the universal embedding concept and SUE, as well as the necessity of the CCA step in the current implementation.
@@ A.11 Ablation Visualization
Appendix B Modality Gap
Recent research involved the modality gap in multimodal learning [49, 38, 65]. That is, the embeddings of the two modalities of contrastive-based methods are located in two completely separate regions of the embedding space. Some efforts have been made to mitigate this gap [39, 23]. In particular, the modality gap means that the embeddings are not universal, as the modalities lay on different intrinsic manifolds. In contrast, a universal embedding should have no modality gap.
As visualized in Fig. 14, SUE indeed does not contain a modality gap. The embedding discovered by SUE is a shared manifold of the two modalities (unlike CLIP, for instance). This arises from two core components of our method: CCA produces whitened outputs, ensuring that the representations for each modality are centered at the origin, which implies zero modality gap by definition; and the MMD objective penalizes distributional discrepancies, further reducing any residual mismatch between modalities.
@@ Appendix C Relative CLIP Score
The qualitative examples in Fig. 2 and Fig. 4a depict the retrieval weakness in reliably measuring semantic coherence across modalities. We address this by presenting a soft variant of Recall@$k$ for the text-image domain, which we name Relative CLIP Score (RCS). CLIP score (CS) is a popular measure of similarity between images [33], but is difficult to interpret. Following this idea, we define the RCS between a pair of an image query and a retrieved text as the CS between them, relative to the CS between the image and its true pair (the RCS between a text query and a retrieved image is defined analogously). This results in a measurement that spans (approximately) between 0 to 1.
@@
#paired Flickr30k
I2T T2I
SUE (ours) 500 0.667 0.638
CLIP $\sim$ 400M 0.971 0.973
Appendix D Theoretical Support
In this section, we cite Thm. 21 from [7] to support the claims made in Sec. 4.
Theorem 1.
(Bérard et al. [7], Thm. 21) Let $h$ be any metric on $\mathcal{M}$ such that $(1-\epsilon)g\leq h\leq(1+\epsilon)$, $\epsilon<\epsilon_{0}$. We assume furthermore that the metrics under consideration have their Ricci curvatures bounded from below by $(n-1)K^{2}$ for some constant $K$. There exists constants $\eta_{g,i,K}(\epsilon),1\leq i\leq N_{0}$, which go to zero with $\epsilon$, such that to any orthogonal basis $\{\psi_{j}\}$ of eigenfunction of $\Delta_{h}$ one can associate an orthonormal basis $\{\phi_{i}\}$ of eigenfunctions of $\Delta_{g}$ satisfying $\lVert\phi_{i}-\psi_{i}\rVert_{\infty}\leq\eta_{h,i,K}(\epsilon)$ for $i\leq N_{0}$, where $\lVert\cdot\rVert_{\infty}$ is the sup-norm.
Intuitively, the result shows that if two embedders $f,g$ of a common manifold $\mathcal{M}$ have bounded distortion and bounded Ricci curvature, then the eigenfunctions of the associated Laplace-Beltrami (or diffusion) operators on $f(\mathcal{M})$ and $g(\mathcal{M})$ are close in the $L_{\infty}$ sense. For further discussion and explanations we refer to [7].
Appendix E Technical Details
E.1 Datasets
Table 12: Datasets’ properties. #unpaired indicates the number of unpaired samples used by SUE; #paired indicates the number of paired samples used by SUE. Model 1 refers to the unimodal encoder for the first modality, and Model 2 corresponds to the second modality. More details on the Unimodal encoders can be found in App. E.2
Dataset Modality 1 Modality 2 Model 1 Model 2 #unpaired #paired
Flickr30k Images Captions DINOv2 SBERT 27794 500
Polyvore Images Captions FashionSigLIP SBERT 17190 500
MSCOCO Images Captions DINOv2 SBERT 8550 100
Edges2Shoes Images Edges (images) DINOv2 VICReg 12690 50
caption-FFHQ Images Captions DINOv2 SBERT 16958 1000
Office31 Images Images VICReg DINOv2 8000 1000
Handwritten [^2] This dataset consists of two handcrafted feature sets derived from images, and therefore can be viewed as a tabular domain. Tabular Tabular NA NA 1900 100
@@ E.1 Datasets
E.2 Unimodal Models
As extensively discussed in Sec. 1 and Sec. 4, our first step is extracting meaningful (unimodal) embeddings that capture semantic similarity, using pre-trained unimodal models. The unimodal models used for the different datasets include: DINOv2 (ViT-B/14 distilled) [58], VICReg [2], SBERT (all-MiniLM-L6-v2) [63], the Marqo-FashionSigLIP image encoder [^3] https://www.marqo.ai/blog/search-model-for-fashion, and the MPMTS Measurements encoder [41]. Tab. 12 details the models used for each dataset. Tab. 13 details the key properties of each unimodal encoder.
Feature extraction details.
For the text modality, we use Sentence-Transformers (e.g., all-MiniLM), which produce fixed-length sentence embeddings via mean pooling as implemented in the library. For the image modality, we extract features using DINOv2, where the [CLS] token serves as the image embedding vector by default.
@@ E.2 Unimodal Models
E.3 Generation and zero-shot
@@
After obtaining the universal image embeddings from the caption-FFHQ dataset using SUE, we encoded the entire dataset of images with a pre-trained GAN inversion model [77] to obtain the latent representation for all images. Next, we trained a converter model to map the universal image embedding of an input image to its corresponding latent. To achieve this, we used an MSE loss to align the converter’s output with the latent obtained from the GAN’s encoder.
@@
For this experiment, we collected random images of 7 object categories from the ImageNet dataset [17]: cat, dog, man, woman, mountain, shore, snow. Each label was converted into the following sentence: "A photo of a []". For each input image, we computed its universal embedding using SUE, which was trained on Flickr30k and then calculated the cosine similarity between the image embedding and the universal embeddings of all label texts. The label for the image was predicted by selecting the text label that had the highest cosine similarity to the image’s universal embedding.
E.4 Cross-domain Classification
@@
For unimodal models, we used VICReg [2] encoder for the Amazon domain and DINOv2 [58] for the DSLR domain. After obtaining universal embeddings from SUE on the test set, we constructed a KNN classifier using 5 neighbors on one domain and evaluated it on the other, as explained in App. A.4.
E.5 Graph Construction
@@
As unimodal models are often trained using cosine similarity (e.g., [63, 58]), we analogously defined $d$ to be the cosine distance:
@@
E.6 Training procedure
For the parametric computation of the SE, we adopted the training process presented in [67], and trained a neural network capable of approximating the SE on the training set. Each modality was trained independently, and upon completion, the resulting parametric maps were used to compute the SE for the test points of each modality. For further details on the training process for this parametric map, we refer the reader to [67].
@@ E.6 Training procedure
E.7 Data split and preparation
@@ E.8 Hyper-parameters
For the numeric SE, we constructed the graph as outlined in Sec. E.5, using $k=100$ for each point. The Laplacian matrix used is the random walk Laplacian, and we selected 10 eigenvectors, which correspond to the output dimension of the SE.
@@
Computing the SE using the parametric approach involves training a neural network. The network architecture consists of an MLP with hidden layers of sizes $4096$, 40964096, and $1024$ for both networks (one per modality). We set the batch size to 4096, and the learning rate to $10^{-4}$ with a decay factor of 0.1, and the training was run for 100 epochs. The optimizer used is Adam, and the learning rate is adjusted using the PyTorch ReduceLROnPlateau scheduler with a patience of 10. We used the same graph construction as used for the numeric SE, outlined in Sec. E.5.
@@
To train the residual network, we employed an MLP architecture with hidden layers of size $128$, 128128, and $128$, incorporating a residual connection from the input to the output. The optimizer used is AdamW, with a learning rate set to $10^{-3}$. The network was trained for 100 epochs.